The length of a rectangle is 4 in. more than twice its width. If the perimeter of the rectangle is 56 in., find the width of the rectangle.
Answer: The width is ______
Show work:

Answer by tutorcecilia(2152)

Formula for the perimeeter of a rectangle:
P=2[(length + width)]
P=56
Length=2w+4
Width=w
.
so,
56=2[(2w+4)+w)]
56/2=2/2(2w+4+w)]
28=3w+4
28-4=3w
24=3w
24/3=3w/3
8=w
.
So,
P=56
Length=2w+4 = 2(8)+4=16+4=20
Width=w=8
.
Plugging everything back into the formula for a perimeter:
P=2[(length + width)]
56=2(20+8)
56=2(28)
56=56 [checks out]